### v2

# EXAONE 4.0: Unified Large Language Models Integrating Non-reasoning and Reasoning Modes

Abstract
This technical report introduces EXAONE 4.0, which integrates a Non-reasoning mode and a Reasoning mode to achieve both the excellent usability of EXAONE 3.5 and the advanced reasoning abilities of EXAONE Deep. To pave the way for the agentic AI era, EXAONE 4.0 incorporates essential features such as agentic tool use, and its multilingual capabilities are extended to support Spanish in addition to English and Korean. The EXAONE 4.0 model series consists of two sizes: a mid-size 32B model optimized for high performance, and a small-size 1.2B model designed for on-device applications. The EXAONE 4.0 demonstrates superior performance compared to open-weight models in its class and remains competitive even against frontier-class models. The models are publicly available for research purposes and can be easily downloaded via https://huggingface.co/LGAI-EXAONE.
1 Introduction
As part of LG AI Research’s EXAONE foundation model series, the EXAONE language models have been developed to support diverse real-world applications through strong instruction-following and reasoning capabilities.
The previous version, EXAONE 3.5 [research2024exaone35serieslarge], focused on real-world usability by strengthening comprehensive instruction-following abilities, while EXAONE Deep [research2025exaonedeepreasoningenhanced] emphasized reasoning performance, particularly in mathematical and coding domains.
With the upcoming era of agentic AI in mind, EXAONE 4.0 introduces agentic tool use—a core capability for this paradigm—and further advances reasoning abilities.
In terms of tool use, the model is developed to enable the integration of various external tools to develop agents or applications. Regarding reasoning performance, the capabilities of EXAONE 4.0 have been improved by leveraging the validated methodologies developed in EXAONE Deep. Notably, EXAONE 4.0 unifies both Non-reasoning mode—enabling rapid thinking and responses—and Reasoning mode—designed for deep thinking and more accurate answers—into a single model, allowing users to experience both modes within one model.
Compared to previous versions of EXAONE, the number of tokens used during pretraining is significantly increased to bolster world knowledge. To further enhancement of expert knowledge, curating training data from specialized domains such as STEM (Science, Technology, Engineering, and Mathematics) fields plays important role on downstream tasks. Furthermore, the extension of maximum context length of the model to 128K tokens enables handling of various tasks based on significantly longer contexts, thereby improving usability. One notable challenge in processing long contexts is the computational burden of attention calculations. To mitigate this, a hybrid architecture combining global attention and local attention is adopted. This approach minimizes performance degradation while reducing computational costs during training and inference.
Moreover, the EXAONE 4.0 officially add Spanish to its supported languages, expanding beyond its previous bilingual support for English and Korean. The development of Spanish language support was designed to minimize any negative impact on English and Korean performance while maintaining the same tokenizer and vocabulary as the previous EXAONE 3.5 and Deep models.
EXAONE 4.0 particularly excels in areas focused on world knowledge and reasoning, especially in mathematical and coding domains. Despite integrating Non-reasoning mode and Reasoning mode, it secures competitive performance in instruction following. The model also shows commendable performance in long context tasks, particularly excelling in document QA (Question Answering) and RAG (Retrieval Augmented Generation) tasks frequently used by real users. Regarding tool use, it achieves a level comparable to competing models, marking the beginning of the fundamental capabilities essential for the upcoming agentic AI era. Additionally, our supported languages now include English, Korean, and the newly supported Spanish. The EXAONE 4.0 model demonstrates competitive performance in both Korean and Spanish across a diverse range of tasks, including expert-level knowledge and mathematical reasoning.
2 Modeling
2.1 Model Configurations
The EXAONE 4.0 model retains a similar structural framework to the EXAONE 3.5 model, but incorporates several key differences in its architecture. Notably, we modify the approach to the attention mechanism. In the previous EXAONE 3.5 model, every layer utilized global attention, whereas the EXAONE 4.0 model employs a hybrid attention mechanism that combines local attention (sliding window attention type) with global attention in a 3:1 ratio, as illustrated in Figure 1.
Figure 1: Visualization of the hybrid attention mechanism when the window size for local attention (sliding window attention) is set to 3. This figure illustrates how context tokens are processed across layers under the hybrid attention mechanism, highlighting the interaction between local and global attention.
Contrary to past findings that models using global attention across all layers performed better, recent studies [gemma2, gemma3, llama4] suggested that utilizing a larger window size (e.g., from 512 to 1,024 or 4,096) and applying global attention to only a minority of layers can still achieve excellent long-context performance. Additionally, it reported that incorporating small amounts of global attention periodically in combinations with heterogeneous structures like Mamba [nvidia2025nemotronhfamilyaccurateefficient, lieber2024jambahybridtransformermambalanguage, minimax2025minimax01scalingfoundationmodels] helped maintain the ability to understand global context.
In designing the EXAONE 4.0 model, a sliding window size of 4K is selected to minimize any adverse effects on short-context performance. Furthermore, the model does not employ Rotary Position embedding [yang2025ropenopeagainnew] for global attention, ensuring that the model does not develop biases towards length and can maintain a global view. For the design of the local attention mechanism, we do not employ the chunked attention strategy. Instead, we adopt sliding window attention, a well‑established form of sparse attention that offers strong theoretical stability. Unlike chunked attention, sliding window attention benefits from wide support in open‑source frameworks, ensuring robust implementation and ease of integration. To prevent performance degradation in short-context areas during long context fine-tuning, the EXAONE 4.0 model employs a careful data selection methodology and a progressive training recipe, effectively balancing efficiency and performance.
Another significant change in the EXAONE 4.0 model is the repositioning of layer normalization (LayerNorm), as shown in Figure 2. According to recent studies [sun2025cursedepthlargelanguage], some layers that do not significantly impact model performance are found mainly in deep layers. This issue is attributed to the Pre-LN transformer architecture [xiong2020layernormalizationtransformerarchitecture], which enhances stability but leads to exponentially increasing variance in outputs as model depth increases. A simple operation to control variance by providing more scaling to outputs as layers deepen was proposed, but we find that the QK-Reorder-LN method [olmo20252olmo2furious, muennighoff2025olmoeopenmixtureofexpertslanguage], which applies LayerNorm after input queries and keys and performs LayerNorm after attention output, yields better performance on downstream tasks despite consuming more computation. The normalization type RMSNorm, which was applied since EXAONE 3.0, is retained in EXAONE 4.0.
Figure 2: Visualization of repositioning layer normalization. The LayerNorm is applied after input queries and keys, and it is performed after attention output again. The type of normalization is RMSNorm.
Finally, the EXAONE 4.0 model series consists of two configurations: 32B and 1.2B. These models share the same vocabulary, which consists primarily of Korean and English tokens in roughly equal proportions, along with a tiny number of multilingual tokens, as detailed in Table 1.
Model size 32B 1.2B
dd_model 5,120 2,048
Number of layers 64 30
Normalization QK-Reorder-LN QK-Reorder-LN
Non-linearity SwiGLU [shazeer2020gluvariantsimprovetransformer] SwiGLU
Feedforward dimension 27,392 4,096
Attention type Hybrid Global
Head type GQA [ainslie2023gqatraininggeneralizedmultiquery] GQA
Number of heads 40 32
Number of KV heads 8 8
Head size 128 64
Max sequence length 131,072 65,536
RoPE theta [su2023roformerenhancedtransformerrotary] 1,000,000 1,000,000
Tokenizer BBPE [wang2020bbpe] BBPE
Vocab size 102,400 102,400
Tied word embedding False True
Knowledge cut-off Nov. 2024 Nov. 2024
Table 1: Configurations of EXAONE 4.0 language models. Key differences from previous versions include a hybrid attention mechanism and modified normalization.
2.2 Pre-training
The amount of data and computational resources used for pretraining in the EXAONE 4.0 models are summarized in Table 2. For the EXAONE 3.5 32B model, 6.5 trillion tokens are used for pretraining. In comparison, the EXAONE 4.0 32B model doubles this amount, utilizing 14 trillion tokens for pretraining. This increase in data is specifically aimed at enhancing the model’s world knowledge. As will be discussed later, this approach yields noticeable improvements in benchmarks that rely on knowledge, such as MMLU-Redux [gema2025mmlu], where the use of more extensive training data has a demonstrable impact on performance.
Model size 32B 1.2B
Size of pretraining data (tokens) 14T 12T
Amount of computation (FLOPs) $2.69\times 10^{24}$ $8.65\times 10^{22}$
Table 2: Pretraining data size and computational resources used for EXAONE 4.0 language models. EXAONE 4.0 utilizes nearly twice the data of its predecessor, EXAONE 3.5.
Furthermore, as recent studies showed that reasoning performance was significantly influenced by the cognitive behavior [gandhi2025cognitivebehaviorsenableselfimproving] acquired from documents seen during pretraining, we perform rigorous data curation during pretraining to enhance post-training performance.
2.3 Context Length Extension
In the EXAONE 4.0 model, the maximum context length is extended to 128K tokens. To achieve this, we undertake a two-stage context length extension process. Initially, a model pretrained with a context length of 4K tokens is firstly extended to 32K tokens. Subsequently, it is further extended to 128K tokens.
The long-context fine-tuning process is meticulously executed, with the Needle In A Haystack (NIAH) test [niah] at each stage to ensure thorough validation of the model’s performance. This iterative refinement continues until comprehensive optimization is achieved and the “green light” signal is consistently observed across all segments, signifying the successful extension of the context length to 128K tokens without compromising the model’s overall performance.
For the 1.2B model, the context length is extended up to 64K tokens, which is approximately twice as long as the typical maximum length of 32K tokens supported by most models in the 1B-parameter range.
2.4 Post-training
In EXAONE 4.0, multiple stages of training is undertaken to enable the model to respond to a variety of user instructions and integrate Non-reasoning and Reasoning models effectively. The training process is primarily organized into three stages: supervised fine-tuning (SFT), reasoning reinforcement learning (RL), and preference learning to integrate Non-reasoning and Reasoning modes as illustrated in Figure 3.
Figure 3: The post-training pipeline of the EXAONE 4.0. The pipeline consists of five stages, which include supervised fine-tuning (SFT), reinforcement learning (RL), and preference learning.
A significant feature of the post-training phase is the large-scale expansion of the SFT data to enhance performance in an efficient manner. To improve reasoning capabilities, RL is employed. Additionally, a hybrid reward mechanism is used in a two-stage preference learning process to seamlessly integrate Non-reasoning and Reasoning modes.
2.4.1 Large-scale Supervised Fine-tuning
The composition of the SFT dataset is divided into non-reasoning and reasoning data. Furthermore, it can be classified into five distinct domains: World Knowledge, Math/Code/Logic, Agentic Tool Use, Long Context, and Multilinguality. Data collection and generation strategies were differentiated for each purpose and domain, and the detailed methodologies are described below.
World Knowledge
For the world knowledge domain, which encompasses a wide range of fields and levels of difficulty, it is essential to enable the distillation of extensive knowledge. Therefore, we filtered problems collected from web sources based on their educational value, prioritizing the use of high-quality data. Among these, we also sample specialized and high-difficulty data to utilize in training for Reasoning mode.
Math, Code, Logic
For the Math, Code, and Logic tasks, the number of unique problems is relatively limited compared to their importance. This is primarily because establishing accurate ground truth is not only essential but also difficult in these domains, thereby limiting our ability to construct as many high-quality problems as desired. Consequently, rather than create unverifiable problems, we train on diverse responses for queries with verifiable answers, and observe that generating multiple responses per unique query is as effective as increasing the diversity or number of unique queries themselves. Furthermore, in the Reasoning mode, responses for Math and Code domains tend to be longer, which increases the risk of degeneration and language inconsistency; thus, careful filtering is applied. Additionally, for the Code domain, we extend our data collection beyond problem-solving to include a software engineering dataset focused on full stack development, created from code corpora.
Long Context
We construct a long-context SFT dataset from web corpora, focusing on tasks that require comprehensive understanding of extended inputs. To train models to identify and reason over dispersed information, we systematically vary both the context length and the location of key content. The dataset also includes instruction-following queries for long-form generation, allowing models to produce coherent and well-structured long outputs. For Korean, we curate long-context data by refining documents such as legal, administrative, and technical texts. These documents are then restructured to accommodate a diverse range of long-context input formats, ensuring variation in structure and content scope.
Agentic Tool Use
To enhance the model’s capability for agentic tool use, we construct datasets focused on both single-turn and multi-turn tasks, leveraging diverse tool lists. Rather than merely creating datasets for single tool calls, we emphasize the construction of more complex, long-horizon tool-calling data. Accordingly, we develop user-agent conversations that incorporate user interaction, execution feedback from the environment, and iterative reasoning, ultimately guiding the agent to achieve the user’s desired goal. These datasets are organized in multi-step and multi-turn formats to better support the learning of agentic tool use.
Multilinguality
To support both Korean and Spanish, we construct datasets that not only target cultural and historical knowledge specific to each language, but also enable the model to engage in fluent, natural conversations with users. We create new instructions in both languages and additionally leveraged translations of selected existing samples as queries. For Korean, in particular, we curate data to address topics relevant to local education and industry experts, ensuring that the model is well-equipped to handle domain-specific queries from Korean users.
Unified Mode Training
In the combined dataset, the Non-Reasoning data primarily consists of diverse tasks, while the Reasoning data is centered on Math and Code domains. Rather than fine-tuning the two modes sequentially, we combine both modes and train them together. The ratio between the two modes is determined by the amount of Reasoning mode data. If the token ratio of Reasoning mode is too high, we observe that the model tends to behave as if it is in Reasoning mode even when Non-Reasoning mode is enabled. Through ablation studies, we set the token ratio of Reasoning to Non-Reasoning data to 1.5:1.
After unified Non-Reasoning/Reasoning mode fine-tuning, to address domain imbalance, we perform a second round of training using high-quality Reasoning data from the Code and Tool Use domains, reusing these samples to further enhance the performance.
2.4.2 Reasoning Reinforcement Learning
To enhance the model’s reasoning capabilities, we conduct online reinforcement learning (RL) following supervised fine-tuning (SFT). Previous studies demonstrated that combining the GRPO (Group Relative Policy Optimization) algorithm [shao2024deepseekmathpushinglimitsmathematical] with verifiable rewards [lambert2025tulu3pushingfrontiers] can effectively improve model performance. To address the limitations of existing GRPO, we propose a new algorithm, named AGAPO (Asymmetric Sampling and Global Advantage Policy Optimization).
Our training dataset encompasses curated data across four categories: mathematics, code, science, and instruction following. To focus training on more informative data samples, we perform accuracy-based filtering by generating eight responses from the SFT model and excluding samples where all eight responses are correct, a pre-filtering step that removes problems that are easy for the model to avoid inefficient training.
The reward function used in RL is tailored for each category. For the mathematics category, a rule-based verifier is used to determine correctness. In the code category, a response is considered correct if its final code block passes all associated test cases. For the science category, a rule-based verifier is first applied; if a response is deemed incorrect, an LLM-judge then performs a more flexible verification. Finally, for the instruction following category, a reward of 1 is assigned if all constraints are satisfied, and 0 otherwise.
For the algorithm design, AGAPO comprehensively improves upon existing methods. Its main features are as follows:
  • Remove Clipped Objective. Previous research has questioned the necessity of PPO(Proximal Policy Optimization) [schulman2017proximalpolicyoptimizationalgorithms]’s clip loss [ahmadian2024basicsrevisitingreinforcestyle] and shown that this clipped objective can degrade performance [minimax2025minimaxm1scalingtesttimecompute] by preventing crucial, low-probability tokens from contributing to gradient updates. These tokens are often associated with reflective behaviors that serve as forks in the reasoning path. AGAPO removes the clipping from PPO and instead uses a standard policy gradient loss. This approach is designed to prevent the dropping of these exploratory tokens, allowing for more substantial policy updates while maintaining training stability.
  • Asymmetric Sampling. Previous works filter out samples where all responses were either correct or incorrect [yu2025dapoopensourcellmreinforcement, he2025skyworkopenreasoner1] because they result in a zero advantage for GRPO. However, as recent work has shown the effectiveness of Negative Sample Reinforcement [zhu2025surprisingeffectivenessnegativereinforcement], AGAPO utilizes an asymmetric sampling method that does not discard samples where all responses are incorrect, thereby including a higher proportion of negative feedback. For these all-incorrect samples, a small negative reward is assigned through the advantage calculation, allowing them to be used to guide the model away from erroneous reasoning paths.
  • Group&Global Advantages. GRPO advantage method does not account for the distribution of the entire batch, which makes it difficult to assign appropriate negative rewards to groups of all-incorrect samples. To improve this, AGAPO calculates the advantage in two stages: group and global. First, at the group level, the advantage is computed using the Leave-One-Out(LOO) method [ahmadian2024basicsrevisitingreinforcestyle] within a response group based on verifiable reward accuracy. Next, normalization is performed across the entire batch (global) to calculate a final advantage that considers the full batch distribution.
  • Sequence Level Cumulative KL. To enhance performance while preserving capabilities learned during the SFT stage, a KL penalty is applied. We adopt the sequence-level cumulative KL [tang2025pitfallskldivergencegradient], as proposed in prior research, to ensure the model receives an appropriate gradient during training.
Objective
The AGAPO objective is defined for a question $q$ sampled from the training distribution $P(Q)$. For each question, the current policy $\pi_{\theta}(\cdot\mid q)$ generates a group of $G$ candidate responses, denoted as $O=\{o_{1},\dots,o_{G}\}$. Each response $o_{i}$ is assigned a verifiable reward $r_{i}\in[0,1]$. The objective function maximizes the following:
$\mathcal{J}_{\mathrm{AGAPO}}(\theta)=\mathbb{E}_{q\sim P(Q),\,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta}(O\mid q)}\Bigl[\frac{1}{G}\sum_{i=1}^{G}\Bigl(A_{\mathrm{global},i}\,\log\pi_{\theta}(o_{i}\mid q)-\beta\,D_{\mathrm{KL}}\!\bigl(\pi_{\theta},\pi_{\mathrm{ref}}\bigr)\Bigr)\Bigr].$ (1)
The global advantage $A_{\mathrm{global},i}$ in the objective is calculated in two stages. First, a leave-one-out (LOO) advantage is computed within each group. This advantage is then normalized across the entire mini-batch of size $K=B\times G$ to yield the final global advantage:
$A_{\mathrm{loo},i}=r_{i}-\frac{1}{G-1}\sum_{j\neq i}r_{j},\quad A_{\mathrm{global},i}=\frac{A_{\mathrm{loo},i}-\operatorname{mean}\!\bigl(\{A_{\mathrm{loo},k}\}_{k}\bigr)}{\operatorname{std}\!\bigl(\{A_{\mathrm{loo},k}\}_{k}\bigr)}.$ (2)
2.4.3 Preference Learning with Hybrid Reward
In the RL stage, we aim to enhance accuracy through verifiable rewards, and do not use human preference. In addition, since the model is specialized for reasoning tasks, we observe a decline in performance in other types of tasks. To overcome these limitations, we introduce an additional preference learning phase.
Preference learning is conducted by directly learning human preferences from chosen and rejected data pairs, akin to the Direct Policy Optimization (DPO) framework [rafailov2024directpreferenceoptimizationlanguage]. We employ SimPER [xiao2025simperminimalistapproachpreference], among various reference-free preference optimization methods for this learning process. The dataset for preference learning is constructed based on on-policy responses [meng2024simposimplepreferenceoptimization, lambert2025tulu3pushingfrontiers] generated by the model after completing the RL phase. For each query, we generate 4 to 16 responses per task, and select chosen and rejected responses based on a hybrid reward combining verifiable reward, preference reward, language consistency reward, and conciseness reward, tailored per task.
The training is conducted separately for two stages. In the first stage, we focus on increasing token efficiency by reducing the generation length while maintaining the performance of the reasoning mode. Therefore, for reasoning-related verifiable training data, we combine the verifiable reward with a conciseness reward to select the shortest response among the correct answers as the chosen option. In the second stage, we employ a combination of preference reward and language consistency reward for human alignment. For the Reasoning Mode data, preference labeling is performed only on the final answer after the reasoning process is complete. Furthermore, to ensure stability during the second stage of training, a portion of the data from the first stage is sampled and reused.
2.5 Data Compliance
Developing AI models requires a large amount of data, and the acquisition and utilization of this data can lead to various legal issues, such as copyright infringement, intellectual property infringement, and personal information protection violations. To minimize these risks, LG AI Research conducts AI Compliance reviews throughout the entire process of data collection, AI model training, and information provision. For more detailed information, please refer to the EXAONE 3.0 Technical Report [research2024exaone3078binstruction] and the LG AI Ethics Principles [lgethics].
3 Evaluation
Mid-size Frontier
EXAONE 4.0 32B  (Reasoning) Phi 4 reasoning-plus Magistral Small-2506 Qwen 3 32B  (Reasoning) Qwen 3 235B  (Reasoning) DeepSeek R1 -0528
Type Hybrid Reasoning Reasoning Hybrid Hybrid Reasoning
# Total Params 32.0 B 14.7 B 23.6 B 32.8 B 235 B 671 B
World Knowledge
MMLU-Redux 92.3 90.8 86.8 90.9 92.7 93.4
MMLU-Pro 81.8 76.0 73.4 80.0 83.0 85.0
GPQA-Diamond 75.4 68.9 68.2 68.4 71.1 81.0
Math / Coding
AIME 2025 85.3 78.0 62.8 72.9 81.5 87.5
HMMT Feb 2025 72.9 53.6 43.5 50.4 62.5 79.4
LiveCodeBench v5 72.6 51.7 55.8 65.7 70.7 75.2
LiveCodeBench v6 66.7 47.1 47.4 60.1 58.9 70.3
Instruction Following
IFEval 83.7 84.9 37.9 85.0 83.4 80.8
Multi-IF  (EN) 73.5 56.1 27.4 73.4 73.4 72.0
Agentic Tool Use
BFCL-v3 63.9 N/A 40.4 70.3 70.8 64.7
Tau-Bench (Airline) 51.5 N/A 38.5 34.5 37.5 53.5
Tau-Bench (Retail) 62.8 N/A 10.2 55.2 58.3 63.9
Multilinguality
KMMLU-Pro  (KO) 67.7 55.8 51.5 61.4 68.1 71.7
KMMLU-Redux  (KO) 72.7 62.7 54.6 67.5 74.5 77.0
KSM  (KO) 87.6 79.8 71.9 82.8 86.2 86.7
MMMLU  (ES) 85.6 84.3 68.9 82.8 86.7 88.2
MATH500  (ES) 95.8 94.2 83.5 94.3 95.1 96.0
Table 3: The main evaluation results of EXAONE 4.0 32B Reasoning mode. Missing entries (N/A, Not Applicable) indicate that the corresponding model does not support the given input length or task. Asterisk () indicates that the scores are from each baseline model’s official technical report, blog or leaderboard.
3.1 Benchmarks
We evaluate EXAONE 4.0 on a diverse set of benchmarks spanning 6 categories: World Knowledge, Math/Coding, Instruction Following, Long Context, Agentic Tool Use, and Multilinguality.
  • World Knowledge We select benchmarks to evaluate the extent of our model’s world knowledge, including MMLU-Redux [gema2025mmlu] and MMLU-Pro [wang2024mmluprorobustchallengingmultitask], a refined and extended version of MMLU [hendrycks2021measuringmassivemultitasklanguage]. Additionally, we utilize GPQA-Diamond [rein2023gpqagraduatelevelgoogleproofqa] to assess the expert-level knowledge in Biology, Physics, and Chemistry.
  • Math/Coding Challenging benchmarks in Math and Coding categories are adopted to evaluate the test-time computational capability of EXAONE 4.0. For Math, we utilize two math Olympiad competitions: AIME 2025 [maa2025aime] and HMMT Feb 2025 [balunovic_srimatharena_2025]. For Coding, LiveCodeBench v5 and v6 [jain2024livecodebenchholisticcontaminationfree] are chosen.
  • Instruction Following To evaluate how well our models understand and align with users’ instructions, we select IFEval [zhou2023instructionfollowingevaluationlargelanguage] and Multi-IF [he2024multiifbenchmarkingllmsmultiturn], the latter being an extension of IFEval to support multi-turn and multilingual scenarios. We use only the English subset of Multi-IF to focus on assessing the multi-turn instruction-following ability on English.
  • Long Context To evaluate the model’s ability to understand and solve tasks requiring long-context comprehension, we adopt three representative benchmarks: HELMET [yen2025helmetevaluatelongcontextlanguage], RULER [hsieh2024rulerwhatsrealcontext], and LongBench [bai2024longbenchbilingualmultitaskbenchmark]. These benchmarks collectively cover both synthetic tasks and real-world scenarios. To maintain a coherent evaluation focused on core long-context capabilities, we exclude the LongCite task from HELMET (see Appendix D for further details).
  • Agentic Tool Use With the advancement of LLM-based agents, numerous benchmarks have emerged to evaluate their tool-use capabilities, among which we focus on the two most widely adopted: BFCL-v3 [patil2025bfcl] and Tau-Bench [yao2024taubenchbenchmarktoolagentuserinteraction]. BFCL-v3 evaluates various aspects of function-calling abilities. Tau-Bench assesses tool calling performance through simulated conversations between a user LLM. We utilize gpt-4.1-2025-04-14 model as the user role.
  • Multilinguality Beyond English, we evaluate our models on two additional languages: Korean and Spanish. For Korean, we use KMMLU-Pro [^0] https://huggingface.co/datasets/LGAI-EXAONE/KMMLU-Pro for measuring practical applicability on professional knowledge and KMMLU-Redux [^1] https://huggingface.co/datasets/LGAI-EXAONE/KMMLU-Redux for assessing real-world expert knowledge instead of KMMLU [son-etal-2025-kmmlu] to ensure benchmark reliability. KMMLU have been reported dataset error and contamination issue between pre-training corpora and task dataset. In addition, we employ Korean School Math (KSM) subset of HRM8K [ko2025understandsolvetranslatebridging] to evaluate a wide range of Korean mathematical knowledge from high-school to Olympiad level. To evaluate the models’ ability to handle long-context Korean inputs, we also include an in-house benchmark, Ko-LongBench (Please refer to Appendix D.4 for details). For Spanish, we adopt the translated version of existing benchmarks. To be specific, we use MMMLU (es)  [^2] https://huggingface.co/datasets/openai/MMMLU and MATH500 [lightman2023lets] (es)  [^3] https://huggingface.co/datasets/bezir/MATH-500-multilingual. Furthermore, we assess translation ability using WMT24++ [deutsch2025wmt24expandinglanguagecoverage], a widely-used translation benchmark. We consider only EN and ES pair, and utilize LLM-as-a-judge 444gpt-4.1-2025-04-14 is used for the judge model. We follow reference-based direct assessment method used in WMT24++ [deutsch2025wmt24expandinglanguagecoverage]. The exact prompt used for judge is in Appendix D.5. to score the translation quality.
Mid-size Frontier
EXAONE 4.0 32B  (Non-  reasoning) Phi 4 Mistral Small-2506 Gemma 3 27B Qwen 3 32B  (Non-  reasoning) Qwen 3 235B  (Non-  reasoning) Llama 4 Maverick DeepSeek V3 -0324
Type Hybrid Non- Reasoning Non- Reasoning Non- Reasoning Hybrid Hybrid Non- Reasoning Non- Reasoning
# Total Params 32.0 B 14.7 B 24.0 B 27.4B 32.8 B 235 B 402 B 671 B
World Knowledge
MMLU-Redux 89.8 88.3 85.9 85.0 85.7 89.2 92.3 92.3
MMLU-Pro 77.6 70.4 69.1 67.5 74.4 77.4 80.5 81.2
GPQA-Diamond 63.7 56.1 46.1 42.4 54.6 62.9 69.8 68.4
Math / Coding
AIME 2025 35.9 17.8 30.2 23.8 20.2 24.7 18.0 50.0
HMMT Feb 2025 21.8 4.0 16.9 10.3 9.8 11.9 7.3 29.2
LiveCodeBench v5 43.3 24.6 25.8 27.5 31.3 35.3 43.4 46.7
LiveCodeBench v6 43.1 27.4 26.9 29.7 28.0 31.4 32.7 44.0
Instruction Following
IFEval 84.8 63.0 77.8 82.6 83.2 83.2 85.4 81.2
Multi-IF  (EN) 71.6 47.7 63.2 72.1 71.9 72.5 77.9 68.3
Long Context
HELMET 58.3 N/A 61.9 58.3 54.5 63.3 13.7 N/A
RULER 88.2 N/A 71.8 66.0 85.6 90.6 2.9 N/A
LongBench v1 48.1 N/A 51.5 51.5 44.2 45.3 34.7 N/A
Agentic Tool Use
BFCL-v3 65.2 N/A 57.7 N/A 63.0 68.0 52.9 63.8
Tau-Bench (Airline) 25.5 N/A 36.1 N/A 16.0 27.0 38.0 40.5
Tau-Bench (Retail) 55.9 N/A 35.5 N/A 47.6 56.5 6.5 68.5
Multilinguality
KMMLU-Pro  (KO) 60.0 44.8 51.0 50.7 58.3 64.4 68.8 67.3
KMMLU-Redux  (KO) 64.8 50.1 53.6 53.3 64.4 71.7 76.9 72.2
KSM  (KO) 59.8 29.1 35.5 36.1 41.3 46.6 40.6 63.5
Ko-LongBench  (KO) 76.9 N/A 55.4 72.0 73.9 74.6 65.6 N/A
MMMLU  (ES) 80.6 81.2 78.4 78.7 82.1 83.7 86.9 86.7
MATH500  (ES) 87.3 78.2 83.4 86.8 84.7 87.2 78.7 89.2
WMT24++  (ES) 90.7 89.3 92.2 93.1 91.4 92.9 92.7 94.3
Table 4: The main evaluation results of EXAONE 4.0 32B Non-reasoning mode. Missing entries (N/A, Not Applicable) indicate that the corresponding model does not support the given input length or task. Asterisk () indicates that the scores are from each baseline model’s official technical report, blog or leaderboard.
3.2 Baselines
To evaluate the performance of language models from various perspectives, recently released open-weight models are selected as baseline models. These baseline models include not only models of similar sizes but also frontier-level models exceeding 100B parameters, which exhibits superior performance. These models can be divided into three types: (1) Non-Reasoning models, which generate their responses in CoT (Chain-of-Thought) style, (2) Reasoning models, which generate in long CoT style, and (3) Hybrid model, which generate in either CoT or long CoT style depending on the mode. Detailed information about the models is presented in the Appendix 8.
3.3 Experimental Setup
Hyperparameters
We sample $n$ different responses for each problem in benchmarks with limited examples to ensure evaluation stability. Specifically, we sample $n=8$ responses for GPQA-Diamond, $n=32$ for AIME 2025 and HMMT Feb 2025, and $n=4$ for LiveCodeBench v5/6, Tau-Bench and MATH500 (es). The accuracy is averaged over the $n$ samples. In Reasoning mode, we set temperature to 0.6, top-p [Holtzman2020The] to 0.95, and apply a presence penalty of 1.5 only for our 32B model. In contrast, for Non-Reasoning mode, greedy decoding is used for a single ($n=1$) generated response, while the same sampling settings as Reasoning mode (except with a presence penalty of 0.0) are used when generating $n>1$ responses. We generate a maximum of 64K tokens for AIME 2025, HMMT Feb 2025, LiveCodeBench v5/6, and KSM benchmarks, while 32K for other benchmarks.
Long-Context Evaluation of Small-Size models
In evaluating long-context performance of Small-Size NON-REASONING models, we extend the context lengths of Qwen3 1.7B and Qwen3 0.6B beyond their 32K token limit by applying YaRN [peng2024yarn], enabling inference up to 64K tokens. For reference, evaluation results of models such as Gemma-3-1B, EXAONE-3.5-2.4B-Instruct, Qwen3 1.7B, and Qwen3 0.6B at context lengths up to 32K tokens are provided in the Appendix D.
Baselines Reproduction
For baseline models, we borrow scores reported in each model’s official technical report, blog, or leaderboard [^5] We refer to https://github.com/LiveCodeBench/submissions for LiveCodeBench, https://matharena.ai/ for HMMT Feb 2025, and https://gorilla.cs.berkeley.edu/leaderboard.html for BFCL-v3. if available. If not, we reproduce the results in our evaluation environment, following the recommended settings when they are explicitly stated [^6] For example, the Qwen3 series explicitly specifies recommended decoding parameters in its Hugging Face repository..
3.4 Experimental Results
Small-size
EXAONE 4.0 1.2B  (Reasoning) EXAONE Deep 2.4B Qwen 3 0.6B (Reasoning) Qwen 3 1.7B  (Reasoning) SmolLM 3 3B  (Reasoning)
Type Hybrid Reasoning Hybrid Hybrid Hybrid
# Total Params 1.28 B 2.41 B 596 M 1.72 B 3.08 B
World Knowledge
MMLU-Redux 71.5 68.9 55.6 73.9 74.8
MMLU-Pro 59.3 56.4 38.3 57.7 57.8
GPQA-Diamond 52.0 54.3 27.9 40.1 41.7
Math / Coding
AIME 2025 45.2 47.9 15.1 36.8 36.7
HMMT Feb 2025 34.0 27.3 7.0 21.8 26.0
LiveCodeBench v5 44.6 47.2 12.3 33.2 27.6
LiveCodeBench v6 45.3 43.1 16.4 29.9 29.1
Instruction Following
IFEval 67.8 71.0 59.2 72.5 71.2
Multi-IF  (EN) 53.9 54.5 37.5 53.5 47.5
Agentic Tool Use
BFCL-v3 52.9 N/A 46.4 56.6 37.1
Tau-Bench (Airline) 20.5 N/A 22.0 31.0 37.0
Tau-Bench (Retail) 28.1 N/A 3.3 6.5 5.4
Multilinguality
KMMLU-Pro  (KO) 42.7 24.6 21.6 38.3 30.5
KMMLU-Redux  (KO) 46.9 25.0 24.5 38.0 33.7
KSM  (KO) 60.6 60.9 22.8 52.9 49.7
MMMLU  (ES) 62.4 51.4 48.8 64.5 64.7
MATH500  (ES) 88.8 84.5 70.6 87.9 87.5
Table 5: The main evaluation results of EXAONE 4.0 1.2B Reasoning mode. Missing entries (N/A, Not Applicable) indicate that the corresponding model does not support the given input length or task. Asterisk () indicates that the scores are from each baseline model’s official technical report, blog or leaderboard.
Small-size
EXAONE 4.0 1.2B  (Non-reasoning) Qwen 3 0.6B (Non-reasoning) Gemma 3 1B Qwen 3 1.7B  (Non-reasoning) SmolLM 3 3B  (Non-reasoning)
Type Hybrid Hybrid Non-Reasoning Hybrid Hybrid
# Total Params 1.28 B 596 M 1.00 B 1.72 B 3.08 B
World Knowledge
MMLU-Redux 66.9 44.6 40.9 63.4 65.0
MMLU-Pro 52.0 26.6 14.7 43.7 43.6
GPQA-Diamond 40.1 22.9 19.2 28.6 35.7
Math / Coding
AIME 2025 23.5   2.6 2.1   9.8   9.3
HMMT Feb 2025 13.0   1.0 1.5 5.1 4.7
LiveCodeBench v5 26.4   3.6 1.8 11.6 11.4
LiveCodeBench v6 30.1   6.9 2.3 16.6 20.6
Instruction Following
IFEval 74.7 54.5 80.2 68.2 76.7
Multi-IF  (en) 62.1 37.5 32.5 51.0 51.9
Long Context
HELMET 41.2 21.1 N/A 33.8 38.6
RULER 77.4 55.1 N/A 65.9 66.3
LongBench v1 36.9 32.4 N/A 41.9 39.9
Agentic Tool Use
BFCL-v3 55.7 44.1 N/A 52.2 47.3
Tau-Bench (Airline) 10.0 31.5 N/A 13.5 38.0
Tau-Bench (Retail) 21.7 5.7 N/A 4.6 6.7
Multilinguality
KMMLU-Pro  (KO) 37.5 24.6 9.7 29.5 27.6
KMMLU-Redux  (KO) 40.4 22.8 19.4 29.8 26.4
KSM  (KO) 26.3 0.1 22.8 16.3 16.1
Ko-LongBench  (KO) 69.8 16.4 N/A 57.1 15.7
MMMLU  (ES) 54.6 39.5 35.9 54.3 55.1
MATH500  (ES) 71.2 38.5 41.2 66.0 62.4
WMT24++  (ES) 65.9 58.2 76.9 76.7 84.0
Table 6: The main evaluation results of EXAONE 4.0 1.2B Non-reasoning mode. Missing entries (N/A, Not Applicable) indicate that the corresponding model does not support the given input length or task. Asterisk () indicates that the scores are from each baseline model’s official technical report, blog, or leaderboard.
Table 345, and 6 present the benchmark performances of our EXAONE 4.0 models in both Reasoning and Non-Reasoning modes. The key results are summarized below:
Superiority in Math/Coding domains
EXAONE 4.0 models demonstrate extraordinary performance in Math/Coding benchmarks. Specifically, EXAONE 4.0 32B model outperforms Qwen3 235B in both Reasoning and Non-Reasoning modes across all Math/Coding benchmarks. At the same time, EXAONE 4.0 1.2B model surpasses all baselines, except for EXAONE Deep 2.4B in Reasoning mode.
Competitive Performance in Tool Use Scenarios
EXAONE 4.0 32B model shows competitive performance in tool use compared to baseline models. For example, in Reasoning mode, it demonstrates similar performance to R1-0528 in Tau-Bench, and achieves comparable BFCL-v3 results with Qwen 3 235B in Non-Reasoning mode. This is noteworthy considering both baselines are much larger than ours. EXAONE 4.0 1.2B model, despite its small size, achieves the highest performance on Tau-Bench (Retail) compared to the baselines.
World Knowledge and GPQA
Both our models excel in benchmarks in the World Knowledge category. Despite their relatively smaller size compared to the baselines, they achieve competitive performance. Among the benchmarks, the EXAONE 4.0 models especially demonstrate better performance in GPQA-Diamond. Both EXAONE 4.0 32B and 1.2B models achieve second-highest performance in GPQA-Diamond when Reasoning mode is available.
3.5 Reasoning Budget
[…]
Reasoning Budget 64K 32K 16K 8K 4K 2K 1K
EXAONE 4.0 32B
AIME 2025 85.3 74.8 44.2 36.8 35.5 35.7 35.6
LiveCodeBench v6 66.7 67.3 53.0 47.6 46.0 45.7 44.0
EXAONE 4.0 1.2B
AIME 2025 45.2 45.3 37.1 24.6 23.2 22.7 22.3
LiveCodeBench v6 45.3 43.0 40.1 38.3 34.0 33.4 29.3
Table 7: The results of controlling the reasoning budget of EXAONE 4.0 models on AIME 2025 and LiveCodeBench v6. The reasoning budget indicates the number of tokens used for reasoning part of the model response. We fix the length of the answer part to 8K.
The result is presented in Table 7. While a reduced reasoning budget leads to some performance degradation, our EXAONE 4.0 models still demonstrate competitive performance even with a 32K reasoning budgets. Specifically, except for the 32B model on AIME 2025, which shows a 12.3% decrease in performance, the decrease for others is similar or less than 5%, maintaining competitive results compared to baseline models.
4 Limitations
EXAONE 4.0 language models, like all existing language models, have certain limitations and may occasionally generate inappropriate responses. The language model generates responses based on the output probability of tokens, and it is determined during learning from training data. While we make every effort to exclude personal, harmful, and biased information from the training data, some problematic content may still be included, potentially leading to undesirable responses. Please note that the text generated by EXAONE 4.0 language models does not reflect the views of LG AI Research.
  • Inappropriate answers may be generated, which contain personal, harmful or other inappropriate information.
  • Biased responses may be generated, which are associated with age, gender, race, and so on.
  • The generated responses rely heavily on statistics from the training data, which can result in the generation of semantically or syntactically incorrect sentences.
  • Since the models do not reflect the latest information, the responses may be false or contradictory.
LG AI Research strives to reduce potential risks that may arise from EXAONE 4.0 language models. Users are not allowed to engage in any malicious activities (e.g., keying in illegal information) that may induce the creation of inappropriate outputs violating LG AI’s ethical principles when using EXAONE 4.0 language models.
5 Deployment
Section B in the Appendix provides license information for using the EXAONE 4.0 models. Understanding the license information is essential for the legal utilization of the language model.
6 Conclusion
In this technical report, we introduce EXAONE 4.0, which integrates non-reasoning mode and reasoning mode. The key features of EXAONE 4.0 include enhancing the practical usability and reasoning capabilities previously supported in EXAONE 3.5 and EXAONE Deep, consolidating them into a single model, and introducing new functionalities such as agentic tool use and support for Spanish. In terms of performance, EXAONE 4.0 demonstrates superior results compared to models of similar scale and achieves competitive performance even compared to frontier models. As part of our future work, we aim to continuously strengthen usability by gradually expanding the supported languages.
Since the release of EXAONE 3.0, LG AI Research has contributed to the expansion of the research ecosystem by publicly disclosing the model in an open-weight format, and has been continuously improving the model based on user feedback. For any improvement suggestions or business-related inquiries regarding the model, please contact us at contact_us@lgresearch.ai.
Appendix A Contributors
All authors are listed in alphabetical order by last name.
Core Contributors
Eunbi Choi, Kibong Choi, Seokhee Hong, Junwon Hwang, Hyojin Jeon, Hyunjik Jo, Jiyeon Jung, Joonkee Kim, Seonghwan Kim, Soyeon Kim, Sunkyoung Kim, Yireun Kim, Yongil Kim, Haeju Lee, Jinsik Lee, Kyungmin Lee, Sangha Park, Heuiyeen Yeen, Hyeongu Yun
Contributors
Kyunghoon Bae, Stanley Jungkyu Choi, Yemuk Choi, Kyubeen Han, Taewan Hwang, Joonwon Jang, Kijeong Jeon, Gerrard Jeongwon Jo, Euisoon Kim, Hyosang Kim, Jihoon Kim, Youchul Kim, Edward Hwayoung Lee, Gwangho Lee, Honglak Lee, Yongmin Park, Young Min Paik, Youngyong Park, Sanghyun Seo, Sihoon Yang, Sihyuk Yi
Appendix B Model License
EXAONE AI Model License Agreement 1.2 - NC
This License Agreement (“Agreement”) is entered into between you (“Licensee”) and LG Management Development Institute Co., Ltd. (“Licensor”), governing the use of the EXAONE AI Model (“Model”). By downloading, installing, copying, or using the Model, you agree to comply with and be bound by the terms of this Agreement. If you do not agree to all the terms, you must not download, install, copy, or use the Model. This Agreement constitutes a binding legal agreement between the Licensee and Licensor.
1. Definitions
1.1 Model: The artificial intelligence model provided by Licensor, which includes any software, algorithms, machine learning models, or related components supplied by Licensor. This definition extends to encompass all updates, enhancements, improvements, bug fixes, patches, or other modifications that may be provided by Licensor from time to time, whether automatically or manually implemented.
1.2 Derivatives: Any modifications, alterations, enhancements, improvements, adaptations, or derivative works of the Model created by Licensee or any third party. This includes changes made to the Model’s architecture, parameters, data processing methods, or any other aspect of the Model that results in a modification of its functionality or output.
1.3 Output: Any data, results, content, predictions, analyses, insights, or other materials generated by the Model or Derivatives, regardless of whether they are in their original form or have been further processed or modified by the Licensee. This includes, but is not limited to, textual or numerical produced directly or indirectly through the use of the Model.
1.4 Licensor: LG Management Development Institute Co., Ltd., the owner, developer, and provider of the EXAONE AI Model. The Licensor holds all rights, title, and interest in the Model and is responsible for granting licenses to use the Model under the terms specified in this Agreement.
1.5 Licensee: The individual, organization, corporation, academic institution, government agency, or other entity using or intending to use the Model under the terms and conditions of this Agreement. The Licensee is responsible for ensuring compliance with the Agreement by all authorized users who access or utilize the Model on behalf of the Licensee.
2. License Grant
2.1 Grant of License: Subject to the terms and conditions outlined in this Agreement, the Licensor hereby grants the Licensee a limited, non-exclusive, non-transferable, worldwide, and revocable license to:
a. Access, download, install, and use the Model solely for research and educational purposes. This includes evaluation, testing, academic research, experimentation, learning, teaching, training and participation in competitions, provided that such participation is in a non-commercial context. Notwithstanding Section 3.1, the Licensee may only provide the Model or Derivatives for a competition if no commercial license is granted to the competition organizer or any third party.
b. Publicly disclose research results and findings derived from the use of the Model or Derivatives, including publishing papers or presentations.
c. Modify the Model and create Derivatives based on the Model, provided that such modifications and Derivatives are used exclusively for research and educational purposes. The Licensee may conduct experiments, perform analyses, and apply custom modifications to the Model to explore its capabilities and performance under various scenarios. If the Model is modified, the modified Model must include “EXAONE” at the beginning of its name.
d. Distribute the Model and Derivatives in each case with a copy of this Agreement.
2.2 Scope of License: The license granted herein does not authorize the Licensee to use the Model for any purpose not explicitly permitted under this Agreement. Any use beyond the scope of this license, including any commercial application or external distribution, is strictly prohibited unless explicitly agreed upon in writing by the Licensor.
3. Restrictions
3.1 Commercial Use: The Licensee is expressly prohibited from using the Model, Derivatives, or Output for any commercial purposes, including but not limited to, developing or deploying products, services, or applications that generate revenue, whether directly or indirectly. Any commercial exploitation of the Model or its derivatives requires a separate commercial license agreement with the Licensor. Furthermore, the Licensee shall not use the Model, Derivatives or Output to develop or improve any models that compete with the Licensor’s models.
3.2 Reverse Engineering: The Licensee shall not decompile, disassemble, reverse engineer, or attempt to derive the source code, underlying ideas, algorithms, or structure of the Model, except to the extent that such activities are expressly permitted by applicable law. Any attempt to bypass or circumvent technological protection measures applied to the Model is strictly prohibited.
3.3 Unlawful Use: The Licensee shall not use the Model and Derivatives for any illegal, fraudulent, or unauthorized activities, nor for any purpose that violates applicable laws or regulations. This includes but is not limited to the creation, distribution, or dissemination of malicious, deceptive, or unlawful content.
3.4 Ethical Use: The Licensee shall ensure that the Model or Derivatives is used in an ethical and responsible manner, adhering to the following guidelines:
a. The Model and Derivatives shall not be used to generate, propagate, or amplify false, misleading, or harmful information, including fake news, misinformation, or disinformation.
b. The Model and Derivatives shall not be employed to create, distribute, or promote content that is discriminatory, harassing, defamatory, abusive, or otherwise offensive to individuals or groups based on race, gender, sexual orientation, religion, nationality, or other protected characteristics.
c. The Model and Derivatives shall not infringe on the rights of others, including intellectual property rights, privacy rights, or any other rights recognized by law. The Licensee shall obtain all necessary permissions and consents before using the Model and Derivatives in a manner that may impact the rights of third parties.
d. The Model and Derivatives shall not be used in a way that causes harm, whether physical, mental, emotional, or financial, to individuals, organizations, or communities. The Licensee shall take all reasonable measures to prevent misuse or abuse of the Model and Derivatives that could result in harm or injury.
4. Ownership
4.1 Intellectual Property: All rights, title, and interest in and to the Model, including any modifications, Derivatives, and associated documentation, are and shall remain the exclusive property of the Licensor. The Licensee acknowledges that this Agreement does not transfer any ownership rights to the Licensee. All trademarks, service marks, and logos associated with the Model are the property of the Licensor.
4.2 Output: Licensor claims no rights in Output. Licensee is solely responsible for the Output and its use.
4.3 Attribution: In any publication or presentation of results obtained using the Model, the Licensee shall provide appropriate attribution to the Licensor, citing the Model’s name and version, along with any relevant documentation or references specified by the Licensor.
5. No Warranty
5.1 “As-Is” Basis: The Model, Derivatives, and Output are provided on an “as-is” and “as-available” basis, without any warranties or representations of any kind, whether express, implied, or statutory. The Licensor disclaims all warranties, including but not limited to, implied warranties of merchantability, fitness for a particular purpose, accuracy, reliability, non-infringement, or any warranty arising from the course of dealing or usage of trade.
5.2 Performance and Reliability: The Licensor does not warrant or guarantee that the Model, Derivatives or Output will meet the Licensee’s requirements, that the operation of the Model, Derivatives or Output will be uninterrupted or error-free, or that defects in the Model will be corrected. The Licensee acknowledges that the use of the Model, Derivatives or Output is at its own risk and that the Model, Derivatives or Output may contain bugs, errors, or other limitations.
5.3 No Endorsement: The Licensor does not endorse, approve, or certify any results, conclusions, or recommendations derived from the use of the Model. The Licensee is solely responsible for evaluating the accuracy, reliability, and suitability of the Model for its intended purposes.
6. Limitation of Liability
6.1 No Liability for Damages: To the fullest extent permitted by applicable law, in no event shall the Licensor be liable for any special, incidental, indirect, consequential, exemplary, or punitive damages, including but not limited to, damages for loss of business profits, business interruption, loss of business information, loss of data, or any other pecuniary or non-pecuniary loss arising out of or in connection with the use or inability to use the Model, Derivatives or any Output, even if the Licensor has been advised of the possibility of such damages.
6.2 Indemnification: The Licensee agrees to indemnify, defend, and hold harmless the Licensor, its affiliates, officers, directors, employees, and agents from and against any claims, liabilities, damages, losses, costs, or expenses (including reasonable attorneys’ fees) arising out of or related to the Licensee’s use of the Model, any Derivatives, or any Output, including any violation of this Agreement or applicable laws.
7. Termination
7.1 Termination by Licensor: The Licensor reserves the right to terminate this Agreement and revoke the Licensee’s rights to use the Model at any time, with or without cause, and without prior notice if the Licensee breaches any of the terms or conditions of this Agreement. Termination shall be effective immediately upon notice.
7.2 Effect of Termination: Upon termination of this Agreement, the Licensee must immediately cease all use of the Model and Derivatives and destroy all copies of the Model and Derivatives in its possession or control, including any backup or archival copies. The Licensee shall certify in writing to the Licensor that such destruction has been completed.
7.3 Survival: The provisions of this Agreement that by their nature should survive termination, including but not limited to, Sections 4 (Ownership), 5 (No Warranty), 6 (Limitation of Liability), and this Section 7 (Termination), shall continue to apply after termination.
8. Governing Law
8.1 Governing Law: This Agreement shall be governed by and construed in accordance with the laws of the Republic of Korea, without regard to its conflict of laws principles.
8.2 Arbitration: Any disputes, controversies, or claims arising out of or relating to this Agreement, including its existence, validity, interpretation, performance, breach, or termination, shall be referred to and finally resolved by arbitration administered by the Korean Commercial Arbitration Board (KCAB) in accordance with the International Arbitration Rules of the Korean Commercial Arbitration Board in force at the time of the commencement of the arbitration. The seat of arbitration shall be Seoul, Republic of Korea. The tribunal shall consist of one arbitrator. The language of the arbitration shall be English.
9. Alterations
9.1 Modifications: The Licensor reserves the right to modify or amend this Agreement at any time, in its sole discretion. Any modifications will be effective upon posting the updated Agreement on the Licensor’s website or through other means of communication. The Licensee is responsible for reviewing the Agreement periodically for changes. Continued use of the Model after any modifications have been made constitutes acceptance of the revised Agreement.
9.2 Entire Agreement: This Agreement constitutes the entire agreement between the Licensee and Licensor concerning the subject matter hereof and supersedes all prior or contemporaneous oral or written agreements, representations, or understandings. Any terms or conditions of any purchase order or other document submitted by the Licensee in connection with the Model that are in addition to, different from, or inconsistent with the terms and conditions of this Agreement are not binding on the Licensor and are void.
By downloading, installing, or using the EXAONE AI Model, the Licensee acknowledges that it has read, understood, and agrees to be bound by the terms and conditions of this Agreement.
Appendix C Baseline models
The models being compared are categorized into open-weight models: Small-size models under 3B, Mid-size models between 10B and 30B, and Frontier models above 200B. Additionally, the models are divided into three types for performance evaluation, with specific details provided in Table 8.
Category Model Parameters Type Release Date
Frontier DeepSeek R1-0528 [deepseekai2025deepseekr1incentivizingreasoningcapability] 671B (MoE) Reasoning May 2025
DeepSeek V3-0324 [deepseekai2025deepseekv3technicalreport] 671B (MoE) Non-reasoning Mar. 2025
Llama 4 Maverick 402B (MoE) Non-reasoning Apr. 2025
Qwen 3 235B [yang2025qwen3technicalreport] 235B (MoE) Hybrid Apr. 2025
Mid-size Qwen 3 32B [yang2025qwen3technicalreport] 32.8B Hybrid Apr. 2025
EXAONE 4.0 32B 32.0B Hybrid Jul. 2025
Gemma 3 27B [gemmateam2025gemma3technicalreport] 27.4B Non-reasoning Mar. 2025
Mistral-Small-3.2-24B-Instruct-2506 24.0B Non-reasoning Jun. 2025
Magistral-Small-2506 [mistralai2025magistral] 23.6B Reasoning Jun. 2025
Phi 4 reasoning plus [abdin2025phi4reasoningtechnicalreport] 14.7B Reasoning Apr. 2025
Phi 4 [abdin2024phi4technicalreport] 14.7B Non-reasoning Dec. 2024
Small-size SmolLM 3 3B 3.08B Hybrid Jul. 2025
EXAONE Deep 2.4B [research2025exaonedeepreasoningenhanced] 2.41B Reasoning Mar. 2025
Qwen 3 1.7B [yang2025qwen3technicalreport] 1.72B Hybrid Apr. 2025
EXAONE 4.0 1.2B 1.28B Hybrid Jul. 2025
Gemma 3 1B [gemmateam2025gemma3technicalreport] 1.00B Non-reasoning Mar. 2025
Qwen 3 0.6B [yang2025qwen3technicalreport] 596M Hybrid Apr. 2025
Table 8: The list of EXAONE 4.0 models and baseline models used for the evaluation along with their parameter size, type, and released date.
Appendix D Evaluation Details
D.1 HELMET
We include the HELMET benchmark [yen2025helmetevaluatelongcontextlanguage] in our evaluation to systematically assess models’ long-context capabilities across both synthetic and real-world tasks. HELMET is designed as a comprehensive suite of diverse, application-centric tasks and addresses key limitations of prior benchmarks, such as inadequate input lengths, over-reliance on retrieval-style setups, and unreliable evaluation metrics. Crucially, it covers a wide spectrum of long-context challenges, including information recall, multi-hop retrieval, in-context generalization, and long-input generation, making it a well-suited benchmark for evaluating models’ ability to process and reason over extended sequences in practical settings.
We adopt six of the seven categories from HELMET, Synthetic Recall, Retrieval-Augmented Generation (RAG), Passage Re-ranking, In-Context Learning (ICL), Long-document Question Answering (LongQA), and Summarization (Summ), to provide a balanced and holistic evaluation of long-context understanding and reasoning abilities.
We formalize our decision to exclude LongCite task from HELMET along three lines:
  • Scope misalignment: HELMET emphasizes general long-context abilities, such as summarization, question answering, retrieval, and reasoning, whereas LongCite centers on sentence-level citation accuracy, which constitutes a distinct attribution task rather than a core comprehension or generative skill.
  • Metric incompatibility: The benchmark employs standardized metrics like SubEM and model-based scoring, while LongCite introduces specialized citation-precision and F1 measures. Integrating these heterogeneous metrics would compromise the uniformity essential for fair model comparison.
  • Benchmark coherence: Including a specialized citation task would divert HELMET from its unified objective of comparing long-context reasoning across models. Such an inclusion would introduce extraneous variability and diminish comparative consistency.
Consequently, omitting LongCite ensures that HELMET remains a concise, cohesive benchmark focused solely on evaluating long-context language modeling capabilities across properly aligned tasks. Detailed task-wise scores are reported in Table 9, while Figure 4 illustrates how performance on each task varies across different context lengths.
Context Len. Model Total Avg. Recall RAG LongQA Summ Rerank ICL
Mid-Size
128K Mistral-Small-2506 61.93 79.82 63.83 70.17 33.32 53.87 70.56
Qwen3 235B 63.33 85.23 63.85 70.26 39.08 52.89 68.68
Qwen3 32B 54.47 74.88 58.20 54.67 36.12 48.17 54.78
Gemma 3 27B 58.34 82.16 64.63 39.59 34.26 55.06 74.38
LlaMA-4-Maverick 13.72 30.00 15.20 14.14 5.12 2.36 15.48
EXAONE 4.0 32B 58.34 94.06 54.75 52.31 25.64 48.78 74.52
Small-Size
32K SmolLM 3B 41.25 75.29 49.75 43.55 18.40 21.92 38.60
Qwen3 1.7B 35.94 50.07 51.00 35.56 17.47 27.89 33.67
Qwen3 0.6B 21.85 40.33 30.50 26.28 12.73 5.51 15.73
Gemma 3 1B 15.49 18.86 32.17 24.70 8.03 2.08 7.13
EXAONE 3.5 2.4B 41.85 73.35 54.08 31.44 18.87 38.69 34.68
EXAONE 4.0 1.2B 42.50 73.52 47.75 30.43 15.23 26.80 61.27
64K SmolLM 3B 38.60 67.81 46.88 41.86 19.58 16.80 38.65
Qwen3 1.7B 33.83 44.29 48.63 36.04 18.68 21.80 33.55
Qwen3 0.6B 21.10 37.13 28.50 27.47 13.85 4.15 15.50
Gemma 3 1B N/A N/A N/A N/A N/A N/A N/A
EXAONE 3.5 2.4B N/A N/A N/A N/A N/A N/A N/A
EXAONE 4.0 1.2B 41.17 69.13 46.25 31.44 15.80 22.33 62.10
Table 9: Comparison of Mid-Size and Small-Size models across tasks on the HELMET benchmark. (N/A) indicates that models not supporting specific input lengths are omitted from evaluation. () denotes that models are extended to 64K context length using YaRN.
Figure 4: Performance of various models across six HELMET task categories, Recall, RAG, Passage Re-ranking, ICL, LongQA, and Summarization, at different context lengths (8K to 128K tokens). Darker cells indicate higher accuracy. Missing entries (N/A) denote models that do not support the corresponding input length or task.
D.2 RULER
We evaluate our model’s long-context capabilities using the RULER benchmark [hsieh2024rulerwhatsrealcontext], a synthetic evaluation suite designed to assess various aspects of long-context understanding beyond simple retrieval. RULER consists of diverse task categories, including retrieval, multi-hop tracing, aggregation, and question answering, and supports flexible configurations for context length and task complexity. The performance of our models across different sequence lengths is summarized in Table 10.
Model 4K 8K 16K 32K 64K 128K
Mid-Size
Mistral-Small-2506 97.18 97.15 96.66 94.57 88.53 71.84
Qwen3 235B 97.70 97.20 96.40 95.10 93.30 90.60
Qwen3 32B 98.40 96.00 96.20 94.40 91.80 85.60
Geamma3 27B 95.5 95.13 93.88 91.1 80.59 66.00
LlaMA-4-Maverick 97.10 96.85 12.96 4.90 4.35 2.85
EXAONE 4.0 32B 96.26 94.85 93.93 93.64 91.73 88.18
Small-Size
SmolLM 3B 92.30 85.01 81.76 77.85 66.27 -
Qwen3 1.7B 89.70 86.58 80.23 75.17 65.94 -
Qwen3 0.6B 80.74 73.64 67.17 60.82 55.09 -
Gemma 3 1B 58.93 46.98 41.09 28.75 N/A -
EXAONE 3.5 2.4B 88.91 87.79 87.27 77.73 N/A -
EXAONE 4.0 1.2B 87.02 86.71 88.83 81.07 77.43 -
Table 10: Accuracy scores of Mid-Size and Small-Size models on the RULER benchmark across varying context lengths (4K to 128K tokens). (N/A) indicates that models not supporting specific input lengths are omitted from evaluation. () denotes that models are extended to 64K context length using YaRN.
D.3 LongBench
LongBench [bai2024longbenchbilingualmultitaskbenchmark] has been suggested as a bilingual benchmark to assess long context comprehension in English and Chinese. We focus on the English subsets, specifically Single-doc QA, Multi-doc QA, Summarization, and Few-shot Learning.
The Single-doc QA task covers datasets such as NarrativeQA [kovcisky2018narrativeqa], Qasper [dasigi-etal-2021-dataset], and MultiFieldQA-EN [bai2024longbenchbilingualmultitaskbenchmark]. For the Multi-doc QA task, we employ benchmarks including HotpotQA [yang-etal-2018-hotpotqa], 2WikiMultihopQA [ho-etal-2020-constructing], and MuSiQue [trivedi-etal-2022-musique]. The Summarization task utilizes datasets like GovReport [huang-etal-2021-efficient], QMSum [zhong-etal-2021-qmsum], and MultiNews [fabbri-etal-2019-multi]. For the Few-shot Learning task, we draw from TREC [li-roth-2002-learning] and TriviaQA [joshi-etal-2017-triviaqa]. All evaluations follow the official protocols and metrics defined in LongBench.
Comprehensive results for each task are shown in Table 11.
Context Len. Model Total Avg. Single-doc QA Multi-doc QA Summarization Few-shot Learning
Mid-Size
128K Mistral-Small-2506 51.48 43.73 52.51 28.82 80.87
Qwen3 235B 45.28 41.45 46.96 25.56 67.13
Qwen3 32B 44.24 41.27 47.97 25.73 62.01
Gemma3 27B 51.54 42.65 54.81 24.45 84.26
LlaMA-4-Maverick 34.71 32.72 24.68 23.84 57.58
EXAONE 4.0 32B 48.12 39.40 48.46 27.34 77.28
Small-Size
32K SmolLM 3B 39.85 33.38 18.26 27.94 79.83
Qwen3 1.7B 41.82 33.61 31.87 26.16 75.63
Qwen3 0.6B 32.72 22.75 20.29 23.11 64.73
Gemma 3 1B 34.91 24.85 24.09 21.41 69.29
EXAONE 3.5 2.4B 42.74 35.03 43.11 20.05 72.75
EXAONE 4.0 1.2B 36.75 30.93 34.98 25.14 55.96
64K SmolLM 3B 39.93 33.53 18.27 28.11 79.83
Qwen3 1.7B 41.92 32.01 32.53 25.95 77.19
Qwen3 0.6B 32.44 22.38 21.40 23.15 62.84
Gemma 3 1B N/A N/A N/A N/A N/A
EXAONE 3.5 2.4B N/A N/A N/A N/A N/A
EXAONE 4.0 1.2B 36.93 31.02 35.09 25.28 56.33
Table 11: Task-wise performance of Mid-Size and Small-Size models on the LongBench benchmark across four task categories: Single-doc QA, Multi-doc QA, Summarization, and Few-shot Learning. Each score represents the average accuracy over the English subset of LongBench at specified context lengths. (N/A) indicates that models not supporting specific input lengths are omitted from evaluation. () denotes that models are extended to 64K context length using YaRN.
D.4 Ko-LongBench
Ko-LongBench is an in-house benchmark developed to evaluate long-context understanding in Korean. It consists of multiple tasks, including Document QA, Story Understanding, Dialogue History Understanding, In-Context Learning, Structured QA, and RAG, allowing for a comprehensive assessment of LLMs’ long-context capabilities in real-world scenarios. A detailed overview of the dataset is provided in Table 12, and representative prompt examples for each task are shown in Figures 5 and 6. Table 13 summarizes the average performance of Small-Size models on Ko-LongBench, reporting scores for both 32K and 64K context lengths.
Category Subtask # Samples Description
SingledocQA / MultidocQA Medical 300 Single- and Multi-Document Question Answering in the Medical Domain
Legal 300 Single- and Multi-Document Question Answering in the Legal Domain
Finance 300 Single- and Multi-Document Question Answering in the Finance Domain
Patent 300 Single- and Multi-Document Question Answering in the Patent Domain
Story Understanding Ordering 66 Evaluation of the Ability to Sequence the Given Story
Mixeing 150 Evaluation of the Ability to Infer the number of Mixed Stories
Long-dialogue History Understanding Wrong chatbot 150 Inferring Inconsistencies with Given Information in Multi-turn Dialogues
Wrong inference 150 Inferring Information that cannot be deduced from Multi-turn Dialogues
Topic classify 150 Evaluating the Ability to classify topics in Multi-turn Dialogues
Long In-context Learning Manual QA 150 Evaluation of Information Extraction Ability based on Product Manuals
Many-Shot 150 Evaluation of Information Extraction ability within a Few-shot Context
Long Structued QA Table QA 300 Table-Based Question Answering : Evaluation of Table Interpretation Skills
RAG Manual QA 150 Single-Document Question on Retrieved Document Context
MultiQA 150 Multi-Document Question based on Retrieved Document Context
Total 2766
Table 12: Descriptions of Ko-LongBench.
Ko-LongBench Example (Long-dialogue History Understanding) 다음 문제에 대해 정답을 고르세요. 당신의 최종 정답은 ABCD 중 하나이고, 정답: 뒤에 와야 합니다. 정답을 고르기 전에 차근차근 생각하고 추론하세요. [Dialogue 0] 안녕하세요. 50대 남성입니다. 반갑습니다. 반갑습니다 저는 20대 여성입니다 그러시군요  혹시 거주하는 곳이 어디인가요? 저는 경상도에 거주하고 있어요  선생님은요? 저는 경기도에 거주하고 있어요  혹시 직업이 있으신가요? 저는 아직까지는 학생입니다. 선생님은 있으신가요? 저는 그냥 일반 직장인이랍니다. 혹시 선생님은 시험기간에 밤을 새시나요? 자주 새요.. 미리 해야하는데 …<중략>
질문 : 위 대화들이 주로 다루고 있는 메인 토픽은 무엇인가? A) 미용과 건강>건강 B) 주거와 생활 C) 개인 및 관계>연애/결혼 D) 여가와 오락>게임 정답: {answer}
Figure 5: Example of Long-dialogue History Understanding (Topic classification) in Ko-LongBench.
Ko-LongBench Example (Long Structued QA) 다음 문제에 대해 정답을 고르세요. 당신의 최종 정답은 ABCD 중 하나이고, 정답: 뒤에 와야 합니다. 정답을 고르기 전에 차근차근 생각하고 추론하세요. 문서 1:인천봉화초등학교 다목적교실 조성공사 시방서(건축) <table><tbody><tr><td>바탕의 종류</td><td>도장 종류</td><td>공법</td></tr><tr><td rowspan=’3’>목재면, 플라스터면, 모르타르면, 콘크리트면</td><td>1종</td><td> …<중략>
질문 : 2022년 경기도청 북부청사 소방시설 점검 및 소방 안전관리 대행 용역에서 본관과 별관의 면적 합계는 전체 면적의 약 몇 퍼센트를 차지하는가? A) 약 50% B) 약 60% C) 약 70% D) 약 80% 정답: {answer}
Figure 6: Example of Long Structued QA (Table QA) in Ko-LongBench.
Model Avg. up to 32K Avg. up to 64K
SmolLM 3B 19.3 15.7
Qwen3 1.7B 62.4 57.1
Qwen3 0.6B 18.6 16.4
Gemma 3 1B 6.3 N/A
EXAONE 3.5 2.4B 57.8 N/A
EXAONE 4.0 1.2B 72.0 69.8
Table 13: Average performance of Small-Size models on Ko-LongBench, a multi-task benchmark designed to evaluate long-context understanding in Korean. The left column reports the average scores across all tasks up to 32K context length, while the right column shows the average scores up to 64K. (N/A) indicates that models not supporting specific input lengths are omitted from evaluation. () denotes that models are extended to 64K context length using YaRN.
D.5 WMT24++
Figure 7 presents the prompt used for LLL-as-a-judge in WMT24++ benchmark. We use the same 0-shot prompt from the official WMT24++ paper.
WMT24++ Judge Prompt "You are a professional judge for evaluating the quality of {src_lang} to {tgt_lang} translations suitable for use in {tgt_region}. Based on the source text, the human-written translation, and machine translation surrounded with triple backticks, your task is to assess the quality of the machine translation on a continuous scale from 0 to 100. A score of 0 means "No meaning preserved," then the scale goes through "Some meaning preserved," to "Most meaning preserved and few grammatical mistakes," up to a score of 100, which means "Perfect meaning and grammar." Your output should only include the score from 0 to 100 without any additional text. {src_lang} text: “‘{src_text}”’ {tgt_lang} human translation: “‘{tgt_text}”’ {tgt_lang} machine translation: “‘{model_text}”’
Figure 7: The judge prompt for evaluating translation quality in WMT24++ benchmark.